OpenStudy (anonymous):

Help lol (I have no idea on how to do this) The sequence a(letter n under the a) = 1000(1.03)^n describes the amount of money in an interest-earning account where n is the number of years passed after the account is open opened. A. $11940.5 B. $1125.51 C. $1159.27 D. $1229.87

OpenStudy (nikato):

do you have more to the question?

OpenStudy (anonymous):

sadly no that's all there is to that

OpenStudy (anonymous):

Do I find n?

OpenStudy (nikato):

oh. is this not a multiple choice question? does this question just have 4 parts?

OpenStudy (anonymous):

well it is a multi choice, but this question is very confusing lol

OpenStudy (nikato):

i dont think you can find any amount unless you are given n, or the year the account is open

OpenStudy (anonymous):

omg xD soo confusing ok can we do another one?

OpenStudy (nikato):

yea. i think we dont have enough info to solve. what's your other problem?

OpenStudy (anonymous):

ok thanks and here ya go: What is the domain of the sequence? 2, 10, 50, 250, 1250 A: 2, 10, 250, 1250 <--- (I think this one but I'm just just used to range and domain on coordinates) B: 2, 3, 4, 5, 6 C: 2, 10, 50, 1250 D: 1, 2, 3, 4, 5

OpenStudy (anonymous):

oops c is 2, 8, 40, 200, 1000

OpenStudy (nikato):

no. domain of sequence are counting numbers: 1,2,3,4,5,... all the other choices are range of sequence, or the terms

OpenStudy (anonymous):

so D?

OpenStudy (anonymous):

ok, how about another one?

OpenStudy (nikato):

yup. and sure

OpenStudy (anonymous):

Write the first five terms of the sentence. \[a _{n} = (-2)^{2}\]

OpenStudy (nikato):

is your question correct?

OpenStudy (anonymous):

oops I put 2 not n sorry

OpenStudy (nikato):

ok. much better. to find first term. plug in 1 into n. to find 2nd, plug in 2 to n to find 3rd, plug in 3 into n. and keep going until 5

OpenStudy (anonymous):

oh wow easier than I Thought lol thanks! :D one more?

OpenStudy (nikato):

sure

OpenStudy (anonymous):

Which recursive rule describes the sequence? 3, 1, 5, -3, 13, ... A: \[a _{1} = 3, a _{n} = -2a_{n-1} +7\] B: \[a _{1} = 3, a _{n} = 7a_{n-1} +2\] C: \[a _{1} = -2, a _{n} = 3a_{n-1} +7\] D: \[a _{1} = -2, a _{n} = 7a_{n-1} +3\]

OpenStudy (nikato):

well you can eliminate C and D because the first term is 3 so a1=3

OpenStudy (nikato):

then i recommend plugging into the equation and see which one fits lets try the first one  and look, its the second term in your sequence so this is probably your asnwer